what am i doing wrong?
¼ = 3 – (2x-1)/(x+2)
¼ * 4(x+2) = 3 – {(2x-1)/(x+2)} *4(x+2)
x + 2 = 3 – (2x – 1)4
x + 2 = 3 – (8x – 4)
x + 2 = 7 – 8x
x + 2 – 2 = 7 – 2 – 8x
x +8x = 7 – 8x +8x
9x = 7
9x/9 = 7/9
x = 7/9

Answer by edjones(8007)

¼ = 3 – (2x-1)/(x+2)
x+2=12(x+2)-4(2x-1) multiply each side by 4(x+2) to eliminate fractions. You didn't multiply 4(x+2) by the 3.
x+2=12x+24-8x+4
-3x=26
x=-26/3 divide each side by -3.
Ed